OpenStudy (anonymous):

Evaluate. For x = 3 4x / 10x^2 4(3) / 10 * (3 * 3) 12 / 10 * 9 12 / 90 = 0.13333 (check if I did ok) Evaluate. For x = 3 4x / 10x^2 4(3) / 10 * (3 * 3) 12 / 10 * 9 12 / 90 = 0.13333 (check if I did ok) @Mathematics

OpenStudy (anonymous):

0.01333

OpenStudy (anonymous):

u should hav to multiply first 10 by 3 and then squaring

OpenStudy (anonymous):

12 / 10 * 3 = 30 * 3 = 90 (like this)

OpenStudy (anonymous):

10*3=30 then square it

OpenStudy (anonymous):

hve u got it?

OpenStudy (anonymous):

Thanks.